OpenStudy (anonymous):

Determine the factors of 4x^2 + 5x − 6.

OpenStudy (anonymous):

\[4x ^{2} + 5x - 6 = 0\] \[4x^{2} + 8x - 3x - 6 = 0 \] \[4x(x + 2) - 3(x+2) = 0\] \[(4x-3)(x+2)=0\] So, \[4x - 3 = 0 \] (or)\[x+2=0\] Therefore, \[x = 3/4, -2\]

OpenStudy (anonymous):

(4x – 2)(x + 3) (4x – 3)(x + 2) (4x – 6)(x + 1) (4x – 1)(x + 6) these r the choices O.o

OpenStudy (anonymous):

Okay. Then, it's the second choice.

OpenStudy (anonymous):

The factors are (4x-3) and (x+2). Sorry, I forgot to mention that.